OpenStudy (anonymous):

Solve for x: -3(x + 3) = -3(x + 1) - 5.

OpenStudy (texaschic101):

do you know how to distribute through the parenthesis ?

OpenStudy (texaschic101):

example : 2(x + 5) = 2*x + 2 * 5 = 2x + 10

OpenStudy (anonymous):

the problem is can't be solved.

OpenStudy (anonymous):

yeah i do but when i worked it out my answer didnt match the answer choices

OpenStudy (texaschic101):

-3(x + 3) = -3(x + 1) - 5 -3x - 9 = -3x - 3 - 5 -3x - 9 = -3x - 8 -3x + 3x = -8 + 9 0 = 1 (incorrect) there is no solution for this problem

OpenStudy (texaschic101):

is that one of your answer choices ?

OpenStudy (anonymous):

ohhh okay

OpenStudy (anonymous):

yes it is

OpenStudy (texaschic101):

cool...if the sides do not equal at the end, then there is no solution

OpenStudy (anonymous):

oh okay now i understand, thanks for helping me

OpenStudy (texaschic101):

anytime :)

OpenStudy (anonymous):

if i have anymore questions will you be able to answer them?

OpenStudy (texaschic101):

if I am logged on...and I will be here for a little while longer. I usually show up at many different times during the day, when I can

OpenStudy (anonymous):

oh okay, but thanks for your help cuz i reallllly needed t.. Im Jayla BTW

OpenStudy (texaschic101):

good to meet you Jayla...I am Kellie

OpenStudy (anonymous):

Nice to meet you too Kellie, ttyl :)